# BMFT: Achieving Fairness via Bias-based Weight Masking Fine-tuning

### v1

@@ 2 Method
Bias in model prediction typically originates from two key sources: (1) the entanglement between noisy, harmful (spurious or irrelevant) features and, useful, core features within the feature extractor; and (2) the incorrect composition of representation in the classification layer, leading to the core features information loss [20]. The main idea of BMFT is to start with debiasing the feature extractor, by fine-tuning only the targeted weights that are disproportionately instigating bias. Subsequently, as the first process may affect how the core features are integrated by the classification layer for prediction, BMFT then fine-tunes a reinitialised classification layer.
@@ 2.2 Bias-Importance-Based Mask Generation
$\mathcal{B}(f_{\theta},\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{a}_{i})=\left|\frac{\sum_{i=1}^{N}f_{\theta}(\mathbf{x}_{i})(1-\mathbf{a}_{i})\mathbf{y}_{i}}{\sum_{i=1}^{N}(1-\mathbf{a}_{i})\mathbf{y}_{i}}-\frac{\sum_{i=1}^{N}f_{\theta}(\mathbf{x}_{i})\mathbf{a}_{i}\mathbf{y}_{i}}{\sum_{i=1}^{N}\mathbf{a}_{i}\mathbf{y}_{i}}\right|,$ (2)
@@
We employ Eq. 3 to quantify the importance of each weight that contributes to the prediction. The importance of each weight for the bias can be computed using the second derivative of the bias influence function (that is, Eq. 2). Noting that the bias influence function is a linear function of the prediction, and thus replaces the loss function with the bias influence function during backpropagation, the importance of a parameter w.r.t. the bias can also be derived by Eq. 3. We denote the weight mask as $M_{i}$, where $i$ is the weight index.
@@ 2.3 Impair-Repair: Two-Step Fine-Tuning
To disentangle core features from biased features, whilst preserving predictive capability, a two-stage fine-tuning process is structured in an “impair-repair” manner. Firstly, we fine-tune the selected masked parameters within the feature extractor $E(\cdot)$ using the group-balanced dataset $\mathcal{D}_{\mathrm{r}}$, guided by an objective function (Eq. 2). Next, we fine-tune the entire reinitialised classification layer $C(\cdot)$. The weights identified by the mask in the feature extractor are closely linked to attribute bias and play a smaller role in predicting the class. This relationship helps us remove the unfavourable influence of bias contained in the model, and yet crucially also retain information embedded in core features.
@@
An unbiased model, trained with a balanced dataset, will have the majority of its classification layer weights being zero, due to minimisation of irrelevant feature impact on class prediction [17]. To this end, we first reinitialise the weights and biases in the classification layer $C(\cdot)$ to zero to enable fast convergence to the optimal unbiased state and then fine-tune the layer on $\mathcal{D}_{\mathrm{r}}$ using a loss which combines the WBCE loss $\mathcal{L}_{\mathrm{WBCE}}$ and fairness constraints (i.e. Eq. 2), as:

### v2

@@ 2 Method
Bias in model prediction typically originates from two key sources: (1) the entanglement between noisy, harmful (spurious or irrelevant) features and useful, core features within the feature extractor; and (2) the incorrect composition of representation in the classification layer, leading to the core features information loss [20]. The main idea of BMFT Fig. 1 is to start with debiasing the feature extractor, by fine-tuning only the targeted weights that are disproportionately instigating bias. Subsequently, as the first process may affect how the core features are integrated by the classification layer for prediction, BMFT then fine-tunes a reinitialised classification layer.
@@ 2.2 Bias-Importance-Based Mask Generation
$\mathcal{B}(f_{\theta},\mathbf{x}_{i},\mathbf{y}_{i},\mathbf{a}_{i})=tpr+fpr,$ (2)
where
[…]
@@
We employ Eq. 3 to quantify the importance of each weight that contributes to the prediction. The importance of each weight for the bias can be computed using the second derivative of the bias influence function (that is, Eq. 2.2). Noting that the bias influence function is a linear function of the prediction, and thus replaces the loss function with the bias influence function during backpropagation, the importance of a parameter w.r.t. the bias can also be derived by Eq. 3. We denote the weight mask as $M_{i}$, where $i$ is the weight index.
@@ 2.3 Impair-Repair: Two-Step Fine-Tuning
To disentangle core features from biased features, whilst preserving predictive capability, a two-stage fine-tuning process is structured in an “impair-repair” manner. Firstly, we fine-tune the selected masked parameters within the feature extractor $E(\cdot)$ using the group-balanced dataset $\mathcal{D}_{\mathrm{r}}$, guided by an objective function (Eq. 2.2). Next, we fine-tune the entire reinitialised classification layer $C(\cdot)$. The weights identified by the mask in the feature extractor are closely linked to attribute bias and play a smaller role in predicting the class. This relationship helps us remove the unfavourable influence of bias contained in the model, and yet crucially also retain information embedded in core features.
@@
An unbiased model, trained with a balanced dataset, will have the majority of its classification layer weights being zero, due to minimisation of irrelevant feature impact on class prediction [17]. To this end, we first reinitialise the weights and biases in the classification layer $C(\cdot)$ to zero to enable fast convergence to the optimal unbiased state and then fine-tune the layer on $\mathcal{D}_{\mathrm{r}}$ using a loss which combines the WBCE loss $\mathcal{L}_{\mathrm{WBCE}}$ and fairness constraints (i.e. Eq. 2.2), as: